if Paul can paint a fence in 2 hours and Fred can paint a the same fence in 3 hours.How long would it take for Fred and Paul to paint the fence together.
Answer by checkley77(12844)

1/2+1/3=1/x
(3+2)/2*3=1/x
5/6=1/x cross multiply
5x=6
x=6/5 hours ans. or 1.2 hours.